HETAV PATEL

Grade 11,

a car is started from rest and accelerated till 30 second.if the distance covered by first 10 second is N1,and distance of the next 10 second is N2 and distance of the last 10 second is N3 then find tthe ratio of N1,N2 and N3 ?

a car is started from rest and accelerated till 30 second.if the distance covered by first 10 second is N1,and distance of the next 10 second is N2 and distance of the last 10 second is N3 then find tthe ratio of N1,N2 and N3 ?

Grade:12

1 Answers

Dhruvit Raithatha
6 years ago
\\\text{At T= 10, } N{_1} = ut + \frac{1}{2}at^2 = 0 + \frac{1}{2}a(10)^2 = 50a \\\\ \text{At T= 20, } N{_2} = ut + \frac{1}{2}at^2 - N{_1} = 0 + \frac{1}{2}a(20)^2 - 50a = 150a \\\\ \text{At T= 30, } N{_3} = ut + \frac{1}{2}at^2 - N{_2} = 0 + \frac{1}{2}a(30)^2 - 100a = 300a \\\\ \\ N{_1}:N{_2}:N{_3} = 50a:150a:300a = 1:3:6